# Common or specific source, features or scores; it is all a matter of information

### v1

@@ 3 An analysis of some arguments against score-based methods in a toy example
In [20], the opposite conclusion was drawn. Therefore, this section is devoted to a critical analysis of their arguments.
@@ 3.1 A toy model
The goal in [20] is to “separate the wheat from the chaff” when calculating likelihood ratios in various ways. To that end, they set up a toy model that allows for relatively simple computations. The setup is such that the computations can be done in several ways: score/feature-based as well as common/specific source. They consider two physical sources $S_{x}$ and $S_{y}$. The goal is to evaluate forensic evidence $e=(e_{x},e_{y})$ coming from these sources with respect to the following two hypotheses:
@@
The toy model in [20] is as follows (cf. their section 2.3). A source $S$ is represented by a real number $\theta_{S}$. In the population of sources, the value of $\theta_{S}$ is normally distributed with mean $\mu$ and variance $\sigma_{D}^{2}$. When we measure on source $S$, we obtain a random perturbation of $\theta_{S}$, modeled by a normal distribution with mean 0 and variance $\sigma_{s}^{2}$ (when the measurement is modeled as a reference measurement) or $\sigma_{u}^{2}$ (when the measurement is modeled as a trace measurement). The distinction between $\sigma_{u}$ and $\sigma_{s}$ allows to take into account that measurements on a reference sample can be performed under more ideal conditions and then tend to give results more closely to the actual value $\theta_{S}$ of the source $S$ they come from.
Next, still following [20, 2.3], a common source scenario consists of two measurements $(e_{u_{1}},e_{u_{2}})$. Both observations involve $\sigma_{u}^{2}$ in their model. The authors write $H_{0,cs}$ and $H_{1,cs}$ for the hypotheses that the measurements are on the same, or different sources, and write the resulting likelihood ratio as
@@
the latter expression being correct since different sources are modeled as having independent parameters. The function $f$ represents the normal density with appropriate parameters.
This is contrasted to a specific source scenario, in which we have measurements $e_{u}$ and $e_{s}$. Here, $e_{u}$ is a trace measurement as before, and $e_{s}$ is a reference type measurement on a known source $S$ with parameter $\mu_{d}$. The authors write $H_{0,ss}$ for the hypothesis that $e_{u}$ has source $S$, and $H_{1,ss}$ for the hypothesis that the source of $e_{u}$ is unknown. However, they tacitly assume knowledge of $\mu_{d}$ in their evaluation of likelihoods ${\mathbb{P}}(e_{u},e_{s}\mid H_{\cdot,ss})$. Indeed, they claim that
@@
However, this is only correct when $e_{u}$ is independent of $e_{s}$, given $H_{\cdot,ss}$. This is so when we condition on $\mu_{d}$, but in that case, the observation $e_{s}$ is redundant, representing a superfluous measurement on a source whose characteristics are already known. In other words, if $\mu_{d}$, the parameter of source $S$, were not conditioned on, the LR would become
@@
thus preventing $e_{s}$ to be be omitted from the conditioning. As such, we see that although the notation does not show this, in the specific source case the hypotheses are similar in notation, but those for the specific source include $\mu_{d}$.
In their section 2.4, the authors set up a comparison between LRs obtained with the common source framework and with the specific source framework, seemingly setting up a comparison where $LR_{CS}$ and $LR_{SS}$ as above are computed on the same data: first, data $(\mu_{d},e_{u},e_{s})$ are generated. Then, the data $(e_{u},e_{s})$ need to be used for a common source evaluation. The authors state that “To calculate the common source likelihood ratio using the data generated under the specific source model, we set $e_{u_{1}}=e_{u}$; $e_{u_{2}}=e_{s};\sigma_{u_{1}}^{2}=\sigma_{u}^{2}$ and $\sigma_{u_{2}}^{2}=\sigma_{s}^{2}$”. Thus, the simulation results seem to be presented as though the same data are evaluated by different models. This is, even, explicitly stated in section 2.2. preceding the toy model where we read, about the difference between common source and specific source: “Each scenario results in different likelihood functions for the same information, and in different interpretations of the results of forensic examination” (our italics).
This however is inaccurate, since for $LR_{SS}$ not only $(e_{u},e_{s})$ are evaluated, but the triples $(e_{u},e_{s},\mu_{d})$ in which $\mu_{d}$ is the parameter of the source that $e_{s}$ comes from, and according to $H_{0,ss}$ also $e_{u}$. Again, $e_{s}$ is redundant in this triple. Thus, the actual comparison is between $(e_{u},\mu_{d})$ and $(e_{u},e_{s})$, both for $H_{0}$ versus $H_{1}$. In order to compare the resulting LRs, one should randomly sample $(e_{u},e_{s},\mu_{d})$ under the two hypotheses, and then compare the LR processing all relevant information $(e_{u},\mu_{d})$ with the situation in which one processes $(e_{u},e_{s})$ only. Instead, the authors choose three different cases: they always take $\mu=10,\sigma_{D}^{2}=10,\sigma_{u}^{2}=2$ and then the cases corresponding to choosing $\mu_{d}=9,\sigma_{s}^{2}=1$, or $\mu_{d}=0,\sigma_{s}^{2}=1$, and finally $\mu_{d}=9,\sigma_{s}^{2}=10^{-5}$. In the last case, almost no differences between $LR_{SS}$ and $LR_{CS}$ are obtained, which is quite reasonable since the small $\sigma_{s}^{2}$ means that $e_{s}$ all but reveals $\mu_{d}$. The differences for the other cases lead the authors to dismiss the common source LRs. As an aside, note that the second case is quite irrelevant, since it represents a source with parameter about three standard deviations away from the mean; such sources will only very rarely be observed.
@@
The simulations in [20] do not allow to infer probabilistic assessments of the distribution of specific-source LRs given a certain specific-source LRs, due to the nature of the simulation. Their simulations compare $LR(\mu_{d},e_{u},e_{s})$ with $LR(e_{u},e_{s})$ for a fixed choice of $\mu_{d}$. However, in order to judge how much these LRs differ in practice, we would need a probability distribution describing what $LR(\mu_{d},e_{u},e_{s})$ may result, given $(e_{u},e_{s})$ or given $LR(e_{u},e_{s})$. That is, we would need to update the probability distribution for the source of $e_{u}$ with the two measurements $(e_{u},e_{s})$ or with $LR(e_{u},e_{s})$, once under each hypothesis, in order to then sample $\mu_{d}$ from the updated probability distribution and calculate $LR(\mu_{d},e_{u},e_{s})$.
In contrast, here the value of $\mu_{d}$ is fixed so that no impression can be obtained from the distribution of specific-source LRs one would obtain given an outcome of the common-source LR. The distributions in the simulations are, therefore, of limited value beyond studying what may happen for very specific parameter choices. For completeness, we carry out the required comparison, even if this is strictly speaking not needed since we already know from the general principles above that the data $(e_{u},\mu_{d})$ will on average be more informative than $(e_{u},e_{s})$, and that always only processing $(e_{u},e_{s})$ is also a perfectly reasonable procedure, if the full evaluation is impossible, for whatever reason. In Figures Figure 3.2 and Figure 3.2, simulation results are plotted with $\sigma_{D}^{2}=10,\sigma_{u}^{2}=2,\sigma_{s}^{2}=1,$ all values in accordance with the first two settings of [20]. However, we let the source value $\mu_{d}$ vary in accordance with the population. We believe that this setting should have been used instead of picking specific values for $\mu_{d}$.
Figure 3.1: $H_{0}$
Figure 3.2: $H_{1}$
A similar flaw in the reasoning occurs when the authors discuss score-based likelihood ratios. Here, the score they consider is the Euclidean distance: for the specific source scenario, we know $\mu_{d}$ and $(\mu_{d}-e_{u})^{2}$, and for the common source scenario we know $(e_{u}-e_{s})^{2}$. Again, the results are presented as a comparison of the same data in two different frameworks, and as representative, whereas the opposite is true: different data are processed for the same hypotheses, and the simulations are by no means representative. The authors, again, conclude against the use of common source LRs: “common source score-based models may be convenient to implement, but are not relevant to most examinations of forensic interest, and do not equal to the weight of forensic evidence.”
Clearly, whereas their simulations results are accurate given their model and parameter choice, these conclusions are not justified.
@@ 3.2 Lack of coherence?
Finally, we comment on another claim in [20]. In case of a specific source score based scenario (SSS), they claim that it is “plagued by a fundamental lack of coherence”. With that, they mean the following. Instead of one source modelled by $E_{y}\sim{\mathcal{N}}(\theta_{y},\sigma_{y}^{2}),$ we assume that we have two sources $A$ and $B$:
$\begin{split}E_{A}&\sim{\mathcal{N}}(\theta_{A},\sigma_{A}^{2}),\\ E_{B}&\sim{\mathcal{N}}(\theta_{B},\sigma_{B}^{2}).\end{split}$ (3.1)
Now two hypotheses $H_{A}$ and $H_{B}$ are considered with respect to trace material $e_{x}$:
  • $H_{A}$: $E_{x}\sim{\mathcal{N}}(\theta_{A},\sigma_{x}^{2})$,
  • $H_{B}:$ $E_{x}\sim{\mathcal{N}}(\theta_{B},\sigma_{x}^{2})$.
In other words, $H_{A}$ states that $e_{x}$ stems from source $A$, while $H_{B}$ assumes source $B$ was the origin of $e_{x}.$
The authors now consider two pieces of evidence. In the first case, they assume the value $\theta_{A}$ is known. Then the evidence is $\theta_{A}$, together with the score $(e_{x}-\theta_{A})^{2}.$ In the second case, we know the value $\theta_{B},$ yielding the score $(e_{x}-\theta_{B})^{2}$. Now [20] notes the fact that in general,
$\begin{split}\frac{{\mathbb{P}}((e_{x}-\theta_{A})^{2}\mid\theta_{A},H_{A})}{{\mathbb{P}}((e_{x}-\theta_{A})^{2}\mid\theta_{A},H_{B})}\neq\frac{{\mathbb{P}}((e_{x}-\theta_{B})^{2}\mid\theta_{B},H_{A})}{{\mathbb{P}}((e_{x}-\theta_{B})^{2}\mid\theta_{B},H_{B})}.\end{split}$ (3.2)
In their view, this is highly problematic because it shows that using $\theta_{A}$ instead of $\theta_{B}$ may yield different results. This is however very natural: using different data may of course lead to a different LR. If $\theta_{A}$ and $\theta_{B}$ would both be taken into account, the order in which this is done does not matter. Indeed,
$\begin{split}\frac{{\mathbb{P}}((e_{x}-\theta_{A})^{2},\theta_{B}\mid H_{A})}{{\mathbb{P}}((e_{x}-\theta_{A})^{2},\theta_{B}\mid H_{B})}=\frac{{\mathbb{P}}((e_{x}-\theta_{B})^{2},\theta_{A}\mid H_{A})}{{\mathbb{P}}((e_{x}-\theta_{B})^{2},\theta_{A}\mid H_{B})}.\end{split}$ (3.3)
On a side note, even if the alleged difference caused by the ordering of $H_{A}$ and $H_{B}$ would have existed, it would have had nothing to do with the score-based nature of SSS. Regardless of scores, using different data will generally lead to a different outcome.
4 Proper scoring rules

### v2

@@ 3 An analysis of some arguments against score-based methods in a toy example
In neumann, the opposite conclusion was drawn. Therefore, this section is devoted to a critical analysis of their arguments.
@@ 3.1 A toy model
The goal in neumann is to “separate the wheat from the chaff” when calculating likelihood ratios in various ways. To that end, they set up a toy model that allows for relatively simple computations. The setup is such that the computations can be done in several ways: score/feature-based as well as common/specific source. They consider two physical sources $S_{x}$ and $S_{y}$. The goal is to evaluate forensic evidence $e=(e_{x},e_{y})$ coming from these sources with respect to the following two hypotheses:
@@
The toy model in neumann is as follows (cf. their section 2.3). Any source is represented by a real number that is the expected value of some hypothetical measurement of (a property of) that source. On the population of sources, this expected value is normally distributed with mean $\mu$ and variance $\sigma_{D}^{2}$. When we measure on a source, we obtain a random perturbation of its expected value. If that expected value is $\mu_{d}$ (this corresponds to $\theta_{S}$ for source $S$ in our previously introduced notation, but here we keep the notation of neumann), the outcomes for measurements on this source are modeled by a normal distribution with mean $\mu_{d}$ and variance $\sigma_{s}^{2}$ (when the measurement is modeled as a reference measurement) or variance $\sigma_{u}^{2}$ (when the measurement is modeled as a trace measurement). The distinction between $\sigma_{u}$ and $\sigma_{s}$ allows to take into account that measurements on a reference sample can be performed under more ideal conditions and then tend to give results more closely to the actual value $\mu_{S}$ of the source $S$ they come from.
Next, still following (neumann, 2.3), a common source scenario consists of two measurements $(e_{u_{1}},e_{u_{2}})$. Both observations involve $\sigma_{u}^{2}$ now, since two traces are modeled. The authors write $H_{0,cs}$ and $H_{1,cs}$ for the hypotheses that the measurements are on the same, or different sources, and write the resulting likelihood ratio as
@@
the latter expression using that different sources are modeled as having independent parameters. The function $f$ represents the normal density with appropriate parameters.
This is contrasted to a specific source scenario, in which we have measurements $e_{u}$ and $e_{s}$. Here, $e_{u}$ is a trace measurement as before, and $e_{s}$ is a reference type measurement on a known source $S$ , i.e., with known parameter $\mu_{d}$. The authors write $H_{0,ss}$ for the hypothesis that $e_{u}$ has source $S$, and $H_{1,ss}$ for the hypothesis that the source of $e_{u}$ is unknown. Next, they claim that
@@
However, this is only correct when $e_{u}$ is independent of $e_{s}$, given $H_{\cdot,ss}$. For this to hold, we must also condition on the source value $\mu_{d}$, otherwise $e_{s}$ is informative for the source. If we do condition on $\mu_{d}$, the observation $e_{s}$ is redundant, representing a superfluous measurement on a source whose characteristic $\mu_{d}$ is already known, leading to the above equality. In other words, if $\mu_{d}$, the parameter of source $S$, were not conditioned on, the LR would become
@@
and $e_{s}$ cannot be omitted from the conditioning. Hence, neumann tacitly assume knowledge of $\mu_{d}$ in their definition of the hypotheses $H_{\cdot,ss}$. Apparently, while the specific source and common source hypotheses are very similar in notation, those for the specific source include the value $\mu_{d}$ of the source of $e_{s}$ whereas the hypotheses for the common source LR do not contain such a value, since in this setup there is no (measured) candidate source for any of the measurements. We will see below that this clarification of the notation is key in understanding the logical errors in the obtained conclusions of neumann.
In their section 2.4, the authors set up a comparison between LRs obtained with the common source framework and with the specific source framework, seemingly setting up a comparison where $LR_{CS}$ and $LR_{SS}$ as above are computed on the same data: first, data $(\mu_{d},e_{u},e_{s})$ are generated. Then, the data $(e_{u},e_{s})$ are used for a common source evaluation. The authors state that “To calculate the common source likelihood ratio using the data generated under the specific source model, we set $e_{u_{1}}=e_{u}$; $e_{u_{2}}=e_{s};\sigma_{u_{1}}^{2}=\sigma_{u}^{2}$ and $\sigma_{u_{2}}^{2}=\sigma_{s}^{2}$”. Thus, the simulation results seem to be presented as though the same data are evaluated by different models. This is, even, explicitly stated in section 2.2. preceding the toy model where we read, about the difference between common source and specific source: “Each scenario results in different likelihood functions for the same information, and in different interpretations of the results of forensic examination” (our italics).
This however is inaccurate, since for $LR_{SS}$ not only $(e_{u},e_{s})$ are evaluated, but the triples $(e_{u},e_{s},\mu_{d})$ in which $\mu_{d}$ is the parameter of the source that $e_{s}$ comes from, and according to $H_{0,ss}$ also $e_{u}$. Again, $e_{s}$ is redundant in this triple. Thus, the actual comparison is between $(e_{u},\mu_{d})$ and $(e_{u},e_{s})$, both for $H_{0}$ versus $H_{1}$. In order to compare the resulting LRs, one should randomly sample $(e_{u},e_{s},\mu_{d})$ under the two hypotheses, and then compare the LR processing all relevant information $(e_{u},\mu_{d})$ with the situation in which one processes $(e_{u},e_{s})$ only. Instead, the authors choose three different cases: they fix $\mu=10,\sigma_{D}^{2}=10,\sigma_{u}^{2}=2$ and then consider the cases corresponding to $\mu_{d}=9,\sigma_{s}^{2}=1$, or $\mu_{d}=0,\sigma_{s}^{2}=1$, and finally $\mu_{d}=9,\sigma_{s}^{2}=10^{-5}$. In the last case, almost no differences between $LR_{SS}$ and $LR_{CS}$ are obtained, which is quite reasonable since the small $\sigma_{s}^{2}$ means that $e_{s}$ all but reveals $\mu_{d}$. The differences for the other cases lead the authors to dismiss the common source LRs. As an aside, note that the second case is quite irrelevant, since it represents a source with parameter about three standard deviations away from the mean; such sources will only very rarely be observed.
@@
The simulations in neumann do not allow to infer probabilistic assessments of the distribution of specific-source LRs given a certain specific-source LRs, due to the nature of the simulation. Their simulations compare $LR(\mu_{d},e_{u},e_{s})$ with $LR(e_{u},e_{s})$ for a fixed choice of $\mu_{d}$. However, in order to judge how much these LRs differ in practice, we would need a probability distribution describing what $LR(\mu_{d},e_{u},e_{s})$ may result, given $(e_{u},e_{s})$ or given $LR(e_{u},e_{s})$. That is, we would need to update the probability distribution for the source of $e_{u}$ with the two measurements via $LR(e_{u},e_{s})$, in order to then sample $\mu_{d}$ from the updated probability distribution and obtain the distribution for $LR(\mu_{d},e_{u},e_{s})$.
In contrast, here the value of $\mu_{d}$ is fixed so that no impression can be obtained from the distribution of specific-source LRs that one would obtain given an outcome of the common-source LR. The distributions in the simulations are, therefore, of limited value beyond studying what may happen for very specific parameter choices. For completeness, we carry out the required comparison, even if this is strictly speaking not needed since we already know from the general principles above that the data $(e_{u},\mu_{d})$ will on average be more informative than $(e_{u},e_{s})$, and that always only processing $(e_{u},e_{s})$ is also a perfectly reasonable procedure, if the full evaluation is impossible, for whatever reason. In Figure 3.1, simulation results are plotted with $\mu=10,\sigma_{D}^{2}=10,\sigma_{u}^{2}=2,\sigma_{s}^{2}=1,$ all values in accordance with the first two settings of neumann. However, we let the source value $\mu_{d}$ vary in accordance with the population. We believe that this setting should have been used instead of picking specific values for $\mu_{d}$. As expected, the specific source information generally leads to better results, but the effect is on average not as dramatic as Figure 1 (b) of neumann suggests.
(a) $H_{0}$ true (same source)
(b) $H_{1}$ true (different sources)
Figure 3.1: Comparison of common source likelihood ratios $LR_{CS}$ and specific source likelihood ratios $LR_{SS}$, obtained with the toy model.
Similar flaws occur in the reasoning leading to Figures 3 and 4. Again, whereas their simulations results are accurate given their model and parameter choice, the conclusions are not justified.
@@ 3.2 Lack of coherence?
Finally, we comment on another claim in neumann. In their Section 3.3, they discuss specific source score based likelihood ratios. Such a LR evaluates as evidence the score $\delta(e_{u},e_{s})$ and $e_{s}$ where $e_{s}$ represents, as before, a control measurement on the putative source. It is not clear from the notation whether the ‘true’ value $\mu_{d}$ of this source is considered to be known as part of the hypothesis. This seems likely in view of it being so for $H_{\cdot,ss}$ discussed before. However, if $\mu_{d}$ is known it is clearly suboptimal to take the score $\delta(e_{u},e_{s})$ instead of $\delta(e_{u},\mu_{d})$, which might suggest that $\mu_{d}$ is not conditioned on here. For the analysis below, whether the true value of the source is conditioned on is not important. Now, they claim that this setup is “plagued by a fundamental lack of coherence”. With that, they mean the following. Suppose that, instead of one known source, we now have two known sources $A$ and $B$ with respective parameters $\mu_{a}$ and $\mu_{b}$, and we still have a measurement $e_{u}$ on a trace. Now two hypotheses $H_{A}$ and $H_{B}$ are considered with respect to $e_{u}$: it either is a measurement of source $A$ or of source $B$, corresponding to the two possibilities
  • $H_{A}$: $e_{u}\sim{\mathcal{N}}(\mu_{a},\sigma_{a}^{2})$,
  • $H_{B}:$ $e_{u}\sim{\mathcal{N}}(\mu_{b},\sigma_{b}^{2})$.
In addition we have measurements $e_{a}$ and $e_{b}$ obtained from $A$ and $B$. From these, one computes the scores $\delta(e_{u},e_{a})$ and $\delta(e_{u},e_{b})$. Now (neumann, (10)) notes the fact that in general,
$\begin{split}\frac{{\mathbb{P}}(\delta(e_{u},e_{a})\mid e_{a},H_{A})}{{\mathbb{P}}(\delta(e_{u},e_{a})\mid e_{a},H_{B})}\neq\frac{{\mathbb{P}}(\delta(e_{u},e_{b})\mid e_{b},H_{A})}{{\mathbb{P}}(\delta(e_{u},e_{b})\mid e_{b},H_{B})}.\end{split}$ (3.1)
They see this as showing that the specific source score based likelihood ratio application “is not coherent in general since it potentially does not support the same proposition depending on which one is considered first”. The above inequality however only expresses that using different data may of course lead to a different LR. Here, either the pair $(\delta(e_{u},e_{a}),e_{a})$ is processed, or the pair $(\delta(e_{u},e_{b}),e_{b})$, but never the full data. Suppose that there is no measurement error, so that $\mu_{a}=e_{a}$ and $\mu_{b}=e_{b}$ are known. Then the above inequality simply says that if we evaluate only the distance from $e_{u}$ to $\mu_{a}$, we get another likelihood ratio than when we only evaluate the distance from $e_{u}$ to $\mu_{b}$, which is perfectly logical. If all data would be taken into account, we would arrive at
$\begin{split}\frac{{\mathbb{P}}(\delta(e_{u},e_{a}),\delta(e_{u},e_{b}),e_{a},e_{b}\mid H_{A})}{{\mathbb{P}}(\delta(e_{u},e_{a}),\delta(e_{u},e_{b}),e_{a},e_{b}\mid H_{B})}=\frac{{\mathbb{P}}(\delta(e_{u},e_{a}),\delta(e_{u},e_{b})\mid e_{a},e_{b},H_{A})}{{\mathbb{P}}(\delta(e_{u},e_{a}),\delta(e_{u},e_{b})\mid e_{a},e_{b},H_{B})}.\end{split}$ (3.2)
4 Discussion and conclusions
The central question of this paper is whether or not additional evidence is somehow beneficial. In Sections 2.1 and 2.2 we showed that this is the case for any cost function, at least in expectation. Even if the cost function’s ‘goal’ is to lie about the truth, then one can tell lies better with more information.
All case-by-case inequalities for cost functions derived in vergeer are contained (but noting that he only treated the $n=2$ case), for cost functions $c_{i,j}$, by Theorem 2.1. For instance, consider the situation in which one compares features with scores. Vergeer showed for this case (and, separately, for others) that on average, using scores leads to higher costs than using all data. These conclusions all directly follow from our results. Indeed, one can interpret the ‘prior’ $\pi$ as the information one has after evaluating the score only. This $\pi$ may be the Bayesian update of an earlier prior upon seeing the score. To take the full features into account in addition to the already known score can in our set-up be interpreted as gathering extra data, and this leads therefore to lower expected costs. The same reasoning goes through for all other examples discussed in vergeer.
This principle also shows that calculating a score-based LR, while less informed than using the full data, is on average still better than using only the prior (3.1.3 in vergeer). This directly contradicts the statement made in neumann that within the Bayesian paradigm, “one cannot use score based likelihood ratios”. We have analysed the arguments from neumann in some detail, as we find it important to explain how the authors of that paper arrived at their conclusion, and why this argument is incorrect.
We would like to add that ‘all data’ is in a sense deceiving, since more knowledge may be possible in the future. In the past, DNA typing was carried out with fewer loci than nowadays. For example, a widespread multiplex was the SGMPlus kit which contained 10 autosomal loci meesterslootenbook. More recent kits contain (about 10 to 15) additional loci, so that calculating a likelihood ratio based on SGMPlus has become a score based likelihood ratio, as it comprises a data reduction relative to the larger sets of loci that are nowadays inspected. This example reveals that what we regard as a full feature method today, may be very well be a partial features method in the future, providing another argument against banning score-based methods.
In fact, every method (pathological examples aside) to calculate a likelihood ratio may in fact be regarded as score-based. One method may of course be more score-based than others, but all continuous measurement methods will usually use some form of simplification (e.g., rounding, binning, cleaning data) and thus lose information that could theoretically be incorporated.
The current paper, like vergeer and neumann, has mostly been concerned with fundamental properties of score-based methods. We hope that this work contributes in clearing score-based likelihood ratios of any fundamental suspicion, and that future research can focus on other important aspects such as calibration ferrer; hannig; ramos2021; rodriguez; ypma.
As a consequence, the question is not whether we should use score based methods, but how they can be used best. What are the conditions needed to justify data reduction, and how good are the decisions made? The answer to these questions depends entirely on the context and the objectives, see also garton. Some score-based methods that use substantial reduction (quite naturally) perform badly enzinger, and may thus not be generally recommended for use in legal contexts.
Next, we come back to the perceived difference between the common source LRs versus the specific source LRs. Suppose we have a trace measurement $e_{u}$ and a source measurement $e_{s}$ of source $S$, and that the question is whether $e_{u}$ also is from $S$ or not. Following neumann we only speak of a specific source LR when the parameters $\theta_{S}$ of the source are known, which can be realized for discrete parameters, but will never really be the case for continuously distributed parameters. Hence, the LR is either calculated based on $(e_{u},e_{s},\theta_{s})$ (and called specific source LR) in which case $\theta_{S}=e_{s}$ or $e_{s}$ is redundant, or on $(e_{u},e_{s})$ (and called common source LR). We see no principle difference between these two approaches here: in both cases, the resulting LR involves (for $H_{1}$) to first probabilistically infer the source parameters (if $\theta_{S}$ is known, this is a trivial step) and then integrating the probability to see $e_{u}$ over the resulting probability distribution for $\theta_{S}$. In both cases, the hypotheses are the same, namely whether the source $S$ is the source of $e_{u}$ or not. We reiterate that the terminology (specific source or common source) only indicates the type of information that is processed, it does not indicate different statistical frameworks.
Similarly, if $e_{u_{1}}$ and $e_{u_{2}}$ are measurements on two traces, we may set up the ‘common source’ hypotheses that they are from the same, versus different, sources. However, we may equally well view this as ‘specific source’ LRs for the hypothesis that $e_{u_{2}}$ comes from $S_{1}$, defined as the source of $e_{u_{1}}$, or not. If additional data $e_{s_{1}}$ on this source are known, then we would ideally evaluate $(e_{u_{1}},e_{u_{2}},e_{s_{1}})$. However, when we process only $(e_{u_{1}},e_{u_{2}})$, the LR that we obtain is still a LR for the hypothesis that $S_{1}$, the source of $e_{u_{1}}$ is also the source of $e_{u_{2}}$, albeit one that has ignored relevant information. In our DNA example, this would amount to testing persons $A$ and $B$ for being siblings, ignoring the DNA profiles of the parents of $A$, hence also considering the possibility that $A$ and $B$ might be siblings with other parents. Nevertheless, we do investigate whether the ‘source’ of $A$ (their parents) are also the source of $B$.
In general, we conclude that the common source LR operates on a smaller set of evidence than the specific source LR, but addresses the same question. Common source LRs quantify evidence for the hypotheses that the source of the first trace is (versus is not) also the source of the second trace, just a specific source LRs do, but based on less information. That does not make them wrong or redundant; while suboptimal, it is preferable to use some evidence than no evidence.
If large posterior odds on the formulated hypotheses are obtained based on a subset of the evidence, say with a score-based LR and/or common source LR, then this means (if one of the hypotheses must be true) that $H_{1}$ is likely correct. If that is so, the evidence not analyzed thus far will yield a LR, when evaluated, that in expectation supports $H_{1}$ also. In any case, the probability that new evidence, when evaluated, will yield a LR of at least $t$ in the direction of the hypothesis that is false, is bounded by $1/t$. It is generally of course not impossible that the evidence, unevaluated so far, can cause a large change in evidential value. The probability of that happening, in view of the bounds for misleading evidence, is therefore mostly determined by the probability that the hypothesis that is not supported by the current LR is nevertheless true. That probability depends on the prior odds and the current LR. For example, suppose a large LR for $A$ and $B$ to be siblings is obtained based on their profiles, so that the posterior odds are a million to one. If the profiles of $A$’s parents would now be included for direct comparison with $B$, we would expect the LR to change dramatically towards unrelatedness only when $A$ and $B$ were, in fact, unrelated after all. But that is unlikely, in view of the odds obtained with the evidence that has been duly processed.
So far, we have argued that processing less data than the full data, by using score-based and/or common source LRs, is not in itself a problematic procedure in the sense that it does lead to a decrease in costs of Bayes decisions. That is to say that if an evaluator has only these data, they are better than nothing. But of course, it is often the case that all data, including the part that was not evaluated, have been observed by the evaluator. Then additional attention is warranted. Sometimes, also without a quantitative statistical model for the unevaluated data, it may nevertheless be possible to make qualitative statements and to recognize the rare cases where the extra information does have a strong impact.
An example of such a situation has long existed for the interpretation of DNA mixture profiles: such a profile consists of a set of peaks whose location reveals the genomic variants (called the alleles) and whose height is a measure for the abundance of that allele post PCR and hence also pre PCR. Initially, models that gave a LR based on the recorded alleles were available, but these models could not process peak height data. Thus, instead of the full data, only a part (i.e., the observed alleles) was processed in the resulting LR, while the forensic analysts and interpretators had access to the whole profile. They would then, visually and qualitatively and not leading to additional quantitative LR assessment, inspect the correspondence between a person of interest (PoI) and the trace profile. This was done to assess whether they would estimate, according to their qualitative knowledge and expertise, that including the peak heights would provide further support for contribution of the PoI, if a LR in favour of that hypothesis had been obtained based on the observed alleles without peak heights. Or, conversely, they would first by visual inspection taking alleles and peak heights into account, have to be convinced of the existence of such support, before proceeding to a computation with the statistical model that ignored the peak heights. Of course, if it would be discovered that the peak heights were inconsistent with contribution, while the computed LR supported it, the computed LR would be overruled. This way, the full data were used, but qualitatively only.
It is of course easy to conceive of peak heights that are inconsistent with the PoI’s contribution (and thus would reduce the LR to zero) even when the profile of a PoI and observed alleles of a trace profile lead to a large LR. However, this is, as in the examples for siblings before, mitigated in practice by the fact that when strong evidence is obtained, this usually means that the posterior probability of contribution is large, and for actual contributors the changes in LRs towards non-contribution are bounded by (2.6). Thus, ignoring the peak heights has primarily the effect of losing evidential strength in the direction of the true hypothesis. This is also why, when a peak-height based model is finally introduced, there is no need to re-visit old cases where strong evidence had been obtained. Only in case where the evidence was weak, this may be worthwhile, time and resources permitting.
Thus, we do not believe that any LR model should be discredited on the basis of the information that it does or does not process. It may in fact be preferable to have a more accurate model for a simplification of the data, than a flawed model for the full data. This can perfectly go together with the evaluator also not being discharged of trying to interpret all relevant data they have access to, even if only qualitatively, and to be convinced of the applicability of the statistical model. This, however, is true for any method, evidence type, and hypotheses.
Appendix